OpenStudy (anonymous):

give the exact value of tan(arccos √3/2 + arcsin -3/5.

OpenStudy (anonymous):

what is that is it like this \(\large \rm tan(cos^{-1}(\sqrt{3}/2)+sin^{-1}(-3/5))\) everything inside tan? or not

OpenStudy (anonymous):

yes how you have it written

OpenStudy (anonymous):

Well, then I would use the sum of tangents formula to start off with \[\tan(a + b) = \frac{ tana + \tan b }{ 1- tana \cdot tanb }\] Alternatively, you could use the division of the sum formulas for sin and cos (since values would cancel out) \[\frac{ \sin(a + b) }{ \cos(a+ b) } = \frac{ (sina)(cosb) + (cosa) (sinb) }{ (cosa) (cosb) - (sina)( sinb) }\]

OpenStudy (anonymous):

you need to ask you self what angle gives \(\sqrt{3}/2\) whe you cosine it also the same for -3/5 what angle gives that ratio

OpenStudy (anonymous):

so arccos=tan a and arcsin=tanb???

OpenStudy (anonymous):

arccos = a, arcsin = b. So you would have: \[\frac{ \tan(\cos^{-1}(\sqrt{3}/2)) + \tan(\sin^{-1}(-3/5)) }{ 1-\tan(\cos^{-1}(\sqrt{3}/2))\cdot \tan(\sin^{-1}(-3/5)) }\] For the cosine inverse portions, you can use your knowledge of what angle gives \(\sqrt{3}/2\) when you take the cosine of it (or check a unit circle), or we could use triangles for all of these values. We need a triangle for the inverse sine portions anyway :P

OpenStudy (anonymous):

yes you need some knowledge about some angles and their sin cos and rational trigonometry

OpenStudy (anonymous):

Having \(tan(sin^{-1}(-3/5))\) can be translated as something along the lines of "Some triangle has a sine valueof -3/5. What is the tangent of this same triangle?" Using tht idea, you can solve the \(tan(sin^{-1}(-3/5)\) portion of your fraction. Unit circle and knowledge of it works out just fine for the other values.

OpenStudy (anonymous):



OpenStudy (anonymous):

4

OpenStudy (anonymous):

Right. So if that missing value is 4, what's the tangent of that triangle?

OpenStudy (anonymous):

using Pythagoras you solve for the missing side and then find the tangent

OpenStudy (anonymous):

correct it is 4

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

Tan is OPPOSITE/ADJACENT!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Bingo. So -3/4 can now replace all of the \(tan(sin^{-1}(-3/5))\) portions of our formula. Now we just need the values for \(tan(cos^{-1}(\sqrt{3}/2))\)

OpenStudy (anonymous):

square root 3

OpenStudy (anonymous):

look here http://www.ambrsoft.com/equations/Trigonometric_files/sinvalues.PNG but you need to know this values, you can't do trig without them

OpenStudy (anonymous):

no square root 3 over 3

OpenStudy (anonymous):

no you need to find the angle whose sin is root3/2

OpenStudy (anonymous):

that is 60 degrees

OpenStudy (anonymous):

\[\tan(\cos^{-1}(\sqrt{3}/2)) = \tan(\frac{ \pi }{ 6 }) = \frac{ 1/2 }{ \sqrt{3}/2 } = \frac{ 1 }{ \sqrt{3} }= \frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

i mean cosine sorry not sin

OpenStudy (anonymous):

so now i am going to replace all of the cos values with square root 3 over 3?

OpenStudy (anonymous):

We now have what we need, so let's plug in the values we found :P \[\frac{ \frac{ \sqrt{3} }{ 3 }- \frac{ 3 }{ 4 } }{ 1- \frac{ \sqrt{3} }{ 3 } \cdot \frac{ -3 }{ 4 }}\] You'll want to simplify this now.

OpenStudy (anonymous):

Able to simplify that ok? :)

OpenStudy (anonymous):

workin on it :) i find the LCD for the top?

OpenStudy (anonymous):

That works. You can actually find the common denominator for all the terms. If the denominators of every term in the entire problem are 1, 3, and 4, what would that common denominator be?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yep. So you can actually multiply everything, top and bottom, by 12. That would give you: \[\frac{ 4\sqrt{3}-9 }{ 12 + 3\sqrt{3} }\] I got rid of the minus sign on bottom, there were two negative signs that could cancel. BUt I multiplied every term by 12 as a common denominator and simplified the fraction substantially. Now, normally, they would want you to rationalize the denominator to remove any square roots. WOuld you kno how to do that?

OpenStudy (anonymous):

ya divide the top and bottom by square root 3

OpenStudy (anonymous):

Well, you need to multiply top and bottom by the conjugate actually. Do you know what the conjugate of the denominator would be?

OpenStudy (anonymous):

12-√3/3?

OpenStudy (anonymous):

I unfortunately have to head out. But once you multiply top and bottom by \(12-3\sqrt{3}\) and simplify, you'll have your final answer. Good luck ^_^

OpenStudy (anonymous):

thanks for ur help!

OpenStudy (anonymous):

NP :3

OpenStudy (anonymous):

\[\tan \left(\sin ^{-1}\left(-\frac{3}{5}\right)+\cos ^{-1}\left(\frac{\sqrt{3}}{2}\right)\right)=\frac{1}{39} \left(25 \sqrt{3}-48\right)=-0.12048 \]